Grade 12Algebra

Construct 3 equations starting with x = 2...........

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

First equation is,
Multiply both side by 6
= 6x = 12 ... [equation 1]
Second equation is,
Subtracting 4 from both side,
=6x–4 = 12 -4
= 6x–4 = 8 ... [equation 2]
Third equation is,
Divide both side by 6
= (6x/6)–(4/6) = (8/6)
= x–(4/6) = (8/6) ... [equation 3]

Thanks